# Search and Society Reimagining Information Access for Radical Futures

### v1

@@ 7 Discussion and conclusion
Despite good intentions. We must be vigilant and reflexively critique our impact, whether under the model of existing fairness and ethics research in IR or under the proposed shift. The call for pluralistic sociotechnical imaginaries must not in this context be confused with uncritical acceptance of all possible futures as equally valid or desirable. Instead, this is a call for critical examination of our community’s existing normative values and future aspirations. This work not only involves explicating our sociotechnical imaginaries but also engaging critically with the history of technology [256] and challenging existing and developing future imaginaries that are counter to the goals of universal emanicipation and justice (e.g.[144]). Above all, we should be wary of any promises of the future that further concentrates wealth and power, or advances any notion of altruism in place of structural change [148].
@@
Desired outcome. Having emphasized the importance of theories of change and ensuring that our research has the desired societal impact and not merely constitute an intellectual exercise, it is only fair that we explicate our own desired outcome of this particular work. We authored this paper because we sincerely believe that information access has a critical role to play in determining our collective futures; and that real change can not be realized by fairness and ethics research happening in silos but only when combined with raising social consciousness, organizing, and movement building. We would consider it a failure if this paper is only cited in future IR papers as a passing remark on social responsibility of IR research. Instead, we hope this work sparks many passionate conversations and debates within the community, and radicalizes us to work on issues of social import in collaboration with other disciplines and civil society. But above all, we hope this paper serves as a clarion call to all IR researchers to reflect on why we do what we do. Personally, we hope that the community continues to build technology not just because we love technology itself, but as an act of radical love for all peoples and the worlds we share.
@@
The author of this paper work at a large technology company in the global north. However, the perspectives presented in this work in many ways directly challenge the western, and specifically big tech’s, view of technology and our collective futures.
Acknowledgements
The author gratefully acknowledges feedback on the early drafts from Michael D. Ekstrand and Ida Larsen-Ledet.

### v4

@@ 5 Conclusion
We must be vigilant and reflexively critique our impact, whether under the model of existing fairness and ethics research in IR or under the proposed shift. The call for pluralistic sociotechnical imaginaries must not in this context be confused with uncritical acceptance of all possible futures as equally valid or desirable. Instead, this is a call for critical examination of our community’s existing normative values and future aspirations. This work not only involves explicating our sociotechnical imaginaries but also engaging critically with the history of technology—e.g.(Merchant, 2023)—and challenging harmful silicon valley ideologies that are counter to the goals of universal emanicipation and justice—e.g.(Gebru and Torres, 2023). Above all, we should be wary of any promises of the future that further concentrates wealth and power, or advances any notion of altruism in place of structural change (Giridharadas, 2019b).
Desired outcomes.
@@
Having emphasized the importance of theories of change and ensuring that our research has the desired societal impact and not merely constitute an intellectual exercise, it is only fair that we explicate our own desired outcome of this particular work. We authored this paper because we sincerely believe that information access has a critical role to play in determining our collective futures; and that real change can not be realized by fairness and ethics research happening in silos but only when combined with raising social consciousness, organizing, and movement building. We would consider it a failure if this paper is only cited in future IR papers as a passing remark on social responsibility of IR research. Instead, we hope this work sparks many passionate conversations and debates within the community, and radicalizes us to work on issues of social import in collaboration with other disciplines and civil society. But above all, we hope this paper serves as a clarion call to all IR researchers to reflect on why we do what we do. Personally, we hope that the community continues to build technology not just because we love technology itself, but as an act of radical love for all peoples and the worlds we share. So, we conclude with one final quote for our readers.
@@
The author of this paper works at a large technology company in the global north. However, the perspectives presented in this work is intended to challenge Big Tech and global north’s view of technology and our collective futures.
Acknowledgments and Disclosure of Funding
The author gratefully acknowledges feedback from Asia Biega, Michael D. Ekstrand, and Ida Larsen-Ledet on the various drafts of this paper. No external funding was received in support of this work.